OpenStudy (anonymous):

Indefinite integral of Sin(2x)^2(cos(2x))^2dx

OpenStudy (amistre64):

integration by parts?

OpenStudy (anonymous):

no need :)

OpenStudy (anonymous):

u sub

OpenStudy (amistre64):

still trying to learn intergrals better... I can derive all day long :) but turning around and going back is killin' me :)

OpenStudy (anonymous):

Yea same here. Deriving is easy for me, going back is whats hurting me

OpenStudy (anonymous):

you can take :\[u = \sin^2 2u\]\[du = 4\cos^2 2u\] so : \[=1/4 \int\limits_{}^{} u du\] = 1/4 [ sin^2 2x/2] I guess :)

OpenStudy (anonymous):

or integration by parts if you want, but that'll be complicated ^_^

OpenStudy (anonymous):

orect me if I'm wrong please.

OpenStudy (anonymous):

correct*

OpenStudy (anonymous):

so, basicaly, I could "reverse" these calculations, if i use the antiderivative, right?

OpenStudy (anonymous):

yep ^_^, by using the antiderivative, you get the original function

OpenStudy (anonymous):

i knew it

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Oh! Hmm is \[ \sin ^{2}2u=\sin (2u)^{2}\]

OpenStudy (anonymous):

why sin^2 2u? it should be sin^2 (2x)/2 I jumped a step ahead lol

OpenStudy (anonymous):

yes, that is right

OpenStudy (anonymous):

becareful, put the brackets all over the thing (sin2u)^2 ^_^

OpenStudy (anonymous):

yes, brackets are LIFESAVERS, i remember how i spent 2 sleeples days trying to find a bug in my program code just to realise a pair of brackets was missing and hence the result was rong

OpenStudy (anonymous):

LOL! I don't like that feeling! >_<

OpenStudy (anonymous):

yeah, it really sucks.... debuging can be bloody annoying

OpenStudy (anonymous):

._. tell me abt it

OpenStudy (anonymous):

haha! im doing this calculus to be a doctor...i honestly dont know when im going to integrate anything while being one though lol

OpenStudy (anonymous):

sorius did you get your answer? ^_^

OpenStudy (anonymous):

lol, mathematics is found everywhere

OpenStudy (anonymous):

working on it right now lol

OpenStudy (anonymous):

you use it to compute the area of something

OpenStudy (anonymous):

or anything really, not just area

OpenStudy (anonymous):

there is alot more :)